A survey of 49 retail stores revealed that the average price of a laptop was $650 with a standard deviation of $50

a) What is the 99% confidence interval to estimate the true cost of the tablet?

b) What sample size would be needed to estimate the true average price of table with an error of + or - 5$ and 95% confidence?
Answer #1

Solution :

Given that,

a) Point estimate = sample mean = \bar x =  $650

sample standard deviation = s = $50

sample size = n = 49

Degrees of freedom = df = n - 1 = 49 - 1 = 48

At 99% confidence level

\alpha = 1 - 99%  

\alpha = 1 - 0.99 =0.01

\alpha/2 = 0.005

Z\alpha/2 = t0.005,48 = 2.682

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.682 * (50 / \sqrt 49)

Margin of error = E = 19.16

The 99% confidence interval estimate of the population mean is,

\bar x  ± E  

= 650 ± 19.16

= ( 630.84, 669.16 )

b) margin of error = E = 5

sample size = n = [t\alpha/2,df* s / E]2

n = [2.682 * 50 / 5 ]2

n = 719.31

Sample size = n = 720